A chemist must mix 16 liters of a 40% acid solution with some of a 70% solution to get a 50% solution.How much of the 70% solution should be used.
Found 2 solutions by mananth, Fakefool5:
Answer by mananth(16946)

percent ---------------- quantity
40 ---------------- 16
70 ---------------- x
50 ---------------- 16+x
....
40*16+70 *x=50(16 +x)
640+70x= 800+50x
70x-50x= 800-640
20x=160
/20
x=8 liters 70 %

Answer by Fakefool5(1)

percent ---------------- quantity
40 ---------------- 16
70 ---------------- x
50 ---------------- 16+x
....
40*16+70 *x=50(16 +x)
640+70x= 800+50x
70x-50x= 800-640
20x=160
/20
x=8 liters 70 %